If disposable income increases from $8000 to $8200 billion and the MPC equals .6, then how much has consumption changed?
1
Expert's answer
2017-02-13T09:19:16-0500
с_1=8000*0,6=4800 billion
с_2=8200*0,6=4920 billion
∆c=c_2-c_1=4920-4800=120 billion